# MFC 5.0: An exascale many-physics flow solver

### v3

@@ 3.2.2 Approximate Riemann solver
The Riemann problem at each finite volume interface is solved using either the Harten–Lax–van Leer (HLL) or the Harten–Lax-van Leer contact (HLLC) approximate Riemann solver [31]. The HLL approximate Riemann solver admits three constant states separated by two discontinuities with wave speeds $S_{\mathrm{L}}$ and $S_{\mathrm{R}}$. The state at the interface between cells $i$ and $i+1$ is
$\bm{q}_{i+1/2,j,k}=\begin{cases}\bm{q}^{\mathrm{(L)}}_{i+1/2,j,k}&S_{\mathrm{L}}\geq 0,\\ \bm{q}^{\mathrm{(HLL)}}&S_{\mathrm{L}}\leq 0\leq S_{\mathrm{R}},\\ \bm{q}^{\mathrm{(R)}}_{i+1/2,j,k}&0\geq S_{\mathrm{R}},\end{cases}$
where the constant start state is
$\bm{q}^{\mathrm{(HLL)}}=\frac{S_{\mathrm{R}}\bm{q}^{\mathrm{(R)}}_{i+1/2,j,k}-S_{\mathrm{L}}\bm{q}^{\mathrm{(L)}}_{i+1/2,j,k}+\bm{F}(\bm{q}^{\mathrm{(L)}}_{i+1/2,j,k})-\bm{F}(\bm{q}^{\mathrm{(R)}}_{i+1/2,j,k})}{S_{\mathrm{R}}-S_{\mathrm{L}}},$
which follows from the integral form of the conservation laws and the consistency condition. The wave speeds $S_{\mathrm{L}}$ and $S_{\mathrm{R}}$ are estimated using the left and right state variables $\bm{q}^{\mathrm{(L)}}_{i+1/2,j,k}$ and $\bm{q}^{\mathrm{(R)}}_{i+1/2,j,k}$. In the subsonic case, $S_{\mathrm{L}}\leq 0\leq S_{\mathrm{R}}$, the HLL flux is given by
$\bm{F}^{\mathrm{(HLL)}}=\bm{F}(\bm{q}^{\mathrm{(R)}}_{i+1/2,j,k})+S_{\mathrm{R}}(\bm{q}^{\mathrm{(HLL)}}-\bm{F}(\bm{q}^{\mathrm{(R)}}_{i+1/2,j,k}))$
which satisfies the Rankine–Hugoniot conditions that ensure the conservation of mass, momentum, and energy across the discontinuity. The flux at the cell interface is then
$\bm{F}^{(x)}_{i+1/2,j,k}=\begin{cases}\bm{F}(\bm{q}^{\mathrm{(L)}}_{i+1/2,j,k})&S_{\mathrm{L}}\geq 0,\\ \bm{F}^{\mathrm{(HLL)}}&S_{\mathrm{L}}\leq 0\leq S_{\mathrm{R}},\\ \bm{F}(\bm{q}^{\mathrm{(R)}}_{i+1/2,j,k})&0\geq S_{\mathrm{R}}\end{cases}$
A shortcoming of the HLL approximate Riemann solver is that it neglects to account for the contact discontinuity in the star region between $S_{\mathrm{L}}$ and $S_{\mathrm{R}}$. As a result, the HLL approximate Riemann solver has more significant numerical dissipation than the HLLC approximate Riemann solver, which accounts for this third contact discontinuity. The HLLC approximate Riemann solver represents the four states as
$\bm{q}_{i+1/2,j,k}=\begin{cases}\bm{q}^{\mathrm{(L)}}_{i+1/2,j,k}&S_{\mathrm{L}}\geq 0,\\ \bm{q}^{(*\mathrm{L})}_{i+1/2,j,k}&S_{\mathrm{L}}\leq 0\leq S_{*},\\ \bm{q}^{(*\mathrm{R})}_{i+1/2,j,k}&S_{*}\leq 0\leq S_{\mathrm{R}},\\ \bm{q}^{\mathrm{(R)}}_{i+1/2,j,k}&0\geq S_{\mathrm{R}},\end{cases},$
where $\bm{q}^{(*\mathrm{L})}$ and $\bm{q}^{(*\mathrm{R})}$ are the states to the left and the right of the contact discontinuity in the star region. Continuity of normal velocity ($u_{*\mathrm{L}}=u_{*\mathrm{R}}=u_{*}$) and pressure ($p_{*\mathrm{L}}=p_{*\mathrm{R}}=p_{*}$) are imposed across the contact discontinuity, $v_{*\mathrm{L}}=v_{\mathrm{L}}$ and $w_{*\mathrm{L}}=w_{\mathrm{L}}$ are imposed across the left discontinuity, and $v_{*\mathrm{R}}=v_{\mathrm{R}}$ and $w_{*\mathrm{R}}=w_{\mathrm{R}}$ are imposed across the right discontinuity to give the fluxes
$\bm{F}^{(*\mathrm{L})}$ $=\bm{F}(\bm{q}^{\mathrm{(L)}}_{i+1/2,j,k})+S_{\mathrm{L}}(\bm{q}^{(*\mathrm{L})}_{i+1/2,j,k}-\bm{q}^{\mathrm{(L)}}_{i+1/2,j,k}),$
$\bm{F}^{(*\mathrm{R})}$ $=\bm{F}(\bm{q}^{\mathrm{(R)}}_{i+1/2,j,k})+S_{\mathrm{R}}(\bm{q}^{(*\mathrm{R})}_{i+1/2,j,k}-\bm{q}^{\mathrm{(R)}}_{i+1/2,j,k}),$
for the star states. The resulting HLLC flux is
$\bm{F}^{\mathrm{(HLLC)}}_{i+1/2,j,k}=\begin{cases}\bm{F}(\bm{q}^{\mathrm{(L)}}_{i+1/2,j,k})&S_{\mathrm{L}}\geq 0,\\ \bm{F}^{(*\mathrm{L})}&S_{\mathrm{L}}\leq 0\leq S_{*},\\ \bm{F}^{(*\mathrm{R})}&S_{*}\leq 0\leq S_{\mathrm{R}},\\ \bm{F}(\bm{q}^{\mathrm{(R)}}_{i+1/2,j,k})&0\geq S_{\mathrm{R}}\end{cases}$
Once computed, the HLL and HLLC fluxes are used to calculate the right-hand side of eq. 10. Identical calculations are performed in the $x$- and $z$-directions with indices $\bm{F}_{i,j+1/2,k}$ and $\bm{F}_{i,j,k+1/2}$.
@@ 3.3 Boundary conditions
Extrapolation and Characteristic [32] boundary conditions are implemented using buffer regions outside the domain to support the stencils of the finite volume reconstructions and provide information outside of the domain for characteristic boundary conditions. Subsonic free stream, supersonic free stream, and wall boundary conditions are implemented to handle various simulation configurations. Simple boundary conditions remain a mainstay of MFC, which were established in Bryngelson et al. [1]. These boundary conditions include symmetry, slip and no-slip walls, prescription of inlet conditions along partial boundary regions, and first-order accurate extrapolation for Neumann-like boundary conditions.
@@ 3.4 Time integration
The conservative variables are advanced in time using high-order total variation diminishing (TVD) and strong stability preserving (SSP) explicit Runge–Kutta time steppers [33]. The third-order accurate version is
$\bm{q}^{(1)}_{i,j,k}$ $=\bm{q}^{n}_{i,j,k}+\Delta t\frac{\partial\bm{q}^{n}_{i,j,k}}{\partial t},$
$\bm{q}^{(2)}_{i,j,k}$ $=\frac{3}{4}\bm{q}^{n}_{i,j,k}+\frac{1}{4}\bm{q}^{(1)}_{i,j,k}+\frac{1}{4}\Delta t\frac{\partial\bm{q}^{(1)}_{i,j,k}}{\partial t},$
$\bm{q}^{n+1}_{i,j,k}$ $=\frac{1}{3}\bm{q}^{n}_{i,j,k}+\frac{2}{3}\bm{q}^{(2)}_{i,j,k}+\frac{2}{3}\Delta t\frac{\partial\bm{q}^{(2)}_{i,j,k}}{\partial t},$
where $\bm{q}^{(1)}_{i,j,k}$ and $\bm{q}^{(2)}_{i,j,k}$ are intermediate steps. First- and second-order accurate time-steppers Runge–Kutta are also available.
@@ 4.1.1 Immersed boundary method
Fluid-solid interactions with rigid bodies are simulated using the Ghost-Cell Immersed Boundary Method with an implementation similar to Tseng and Ferziger [34]. This method imposes Neumann boundary conditions for the pressure, densities, and volume fractions and a no-slip or slip boundary condition for the velocity on the immersed boundary.
In MFC 5.0, the solid geometries are parsed. Each cell is characterized either as within a fluid or a solid region. For each immersed boundary, the level set field $\bm{\varphi}$ is calculated. This field stores the vector from each cell to the closest point on the immersed boundary. The ghost points, image points, and interpolation coefficients are calculated and used to enforce the appropriate boundary conditions. These are computed prior to simulation execution and follow the below algorithm.
The algorithm that applies the boundary conditions follows. MFC first parses the STL or OBJ geometry file (or level set). We determine if each finite volume is in a solid, immersed region via its surface normals. The level set field is computed for each immersed boundary. This field stores the vector from each finite volume center to the closest point on the immersed boundary. If each cell center in the solid regions is within three cells of a fluid cell, we characterize it as a ghost point. For each ghost point, we compute the position of its corresponding image point as $\bm{x}_{\mathrm{(ip)}}=\bm{x}_{\mathrm{(gp)}}+2\bm{\varphi}(\bm{x}_{\mathrm{(gp)}})$, where $\bm{x}_{\mathrm{(ip)}}$ is the image point position, $\bm{x}_{\mathrm{(gp)}}$ is the ghost point position, and $\bm{\varphi}$ is the level set field. The fluid properties at the image points are interpolated from surrounding grid cells, so we compute the interpolation coefficients for each image point using second-order accurate inverse distance weighting. Finite volumes in an immersed solid region are excluded from interpolation.
At each step in a flow simulation, we interpolate fluid properties at the image point. For each fluid property $q$ with a Neumann boundary condition, we compute $q_{\mathrm{(ip)}}=q_{\mathrm{(gp)}}$. We set the velocity based on its boundary condition:
$\bm{u}_{\mathrm{(gp)}}=\begin{cases}\bm{0},&\text{no slip,}\\ \bm{u}_{\mathrm{(ip)}}-(\hat{\bm{n}}\cdot\bm{u}_{\mathrm{(ip)}})\hat{\bm{n}},&\text{slip}.\end{cases}$ (11)
@@ 4.1.2 Treatment of complex geometries
We use a ray-tracing algorithm to translate complex geometries in ASCII and stereolithography (STL) format onto computational grids compatible with the immersed boundary method. The ASCII STL format represents geometric models using a triangular mesh. Each model surface is decomposed into a set of non-overlapping triangular facets. Three vertices and a normal vector define each triangular facet. Unlike simple geometries such as cubes and spheres, most complex geometric models do not have analytical solutions for their boundary normal vectors. Consequently, the level sets of complex geometries mentioned in section 4.1.1 also lack analytical solutions. We need to approximate the geometry’s boundary and boundary normal vectors using the STL vertices located on the boundary. We use an edge-manifoldness algorithm based on Weiler [35] to group the boundary STL vertices that are used to determine the level sets and image points. Coarse STL files are interpolated during ray tracing to achieve results consistent with those obtained from higher-resolution STL files. The implementation is validated for a Mach 2 helium flow over a sphere with initial flow density $\rho_{0}$ and $\mbox{\text{Re}}=6.5\times 10^{5}$ in fig. 1.
@@ 4.1.3 Phase change: First order transition, liquid–vapor
The phase change model for $N$ fluids is integrated into MFC implemented by introducing disequilibrium terms for pressure ($p$), temperature ($T$), and chemical potential ($\mu$) into $\bm{s}(\bm{q})$ in 9. This inclusion increases system stiffness, making a fractional-step method essential [36, 37, 4]. In the case of infinitely fast relaxation, the governing PDEs can be replaced with nonlinear, lower-dimensional, algebraic relations. For example, for the $pT$-relaxation, we have [38]
@@
with $\alpha_{m}\rho_{m}=\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2}$ and $e_{*,i}$ an energy reference parameter added to the stiffened gas EOS (see section 2.2) internal energy equation for species $i$. Subscripts $(\,\cdot\,)_{1,2}$ indicate the reacting liquid and vapor. The remaining subscripts are the inert phases. $A$, $B$, $C$, and $D$ are fluid-dependent parameters [4]. These strategies are compatible with five- and six-equation models and are solved via Newton’s method. Once the solution is obtained, the solver advances to the next time step.
@@ 4.1.4 Euler–Euler sub-grid bubble dynamics
We use Euler–Euler sub-grid models for bubble dynamics, including the method of classes and the method of moments. The method of classes is based on the ensemble phase-averaged equations [39], which can be used to represent dispersions of radially oscillating bubbles as well as other particles [40, 41, 42, 43]. The void fraction of the dispersed phase $\alpha$ is assumed to be negligible compared to the liquid phase ($\alpha_{l}$) under dilute assumptions. The bubbles are represented statistically via random variables $R$, $\dot{R}$, and $R_{o}$, corresponding to the instantaneous bubble radius, time derivative, and equilibrium bubble radius. The mixture-averaged pressure field is
@@
where $p_{bw}$ is the associated bubble wall pressure and $p_{\ell}(\bm{x},t)$ is the liquid pressure according to the modified stiffened-gas EOS [44]:
@@
the right-hand side represents the change in void fraction due to bubble growth and collapse. The over-barred terms appearing in the above equations denote average quantities of the bubble dispersion. These are averaged over $N_{\mathrm{bin}}$ bins of $R_{o}$ using a log-normal probability distribution function (pdf). We implement the Rayleigh–Plesset and Keller–Miksis equations [45] for the radial bubble dynamics. The Keller–Miksis dynamics are
@@
The method of classes incorporates stochasticity in $R_{o}$ using a log-normal pdf (probability density function). However, modeling complex bubbly flow phenomena requires the expansion of the set of stochastic variables, including instantaneous bubble variables. Moreover, the inaccuracy of the polytropic assumption can limit the scope of its applications. The method of moments is used for this purpose, which employs a population balance equation (PBE) to govern the pdf of the bubbles. The pdf $f(R,\dot{R},R_{o})$ governs the moments $\mu_{lmn}$ as
@@
The moments are evaluated using a conditional inversion procedure that follows Fox et al. [46]. The conditional probability density function $f(R,\dot{R}|R_{o})$ without coalescence or breakup effects is governed by the PBE
@@
and $\Omega=\Omega_{R}\times\Omega_{\dot{R}}\times\Omega_{R_{o}}=(0,\infty)\times(-\infty,\infty)\times(0,\infty)$ [42]. The integral is computed via quadrature nodes obtained by the inversion procedure.
@@
where $N_{R_{o}}$ represents the polydispersity of the bubble population. The conditional moments $\mu_{lm}$ are evaluated using a conditional hyperbolic inversion procedure (CHyQMOM). The CHyQMOM algorithm is described in Fox et al. [46], and a detailed discussion of its application to bubble dynamics is in Bryngelson et al. [41]. The total moments follow as
@@
These moments evaluate the right-hand side of the moment evolution equation and the ensemble-phase averaged terms. The evaluation of $\overline{R^{3}p_{bw}}$ requires the accounting of statistics of $p_{b}$. Without polytropic assumptions, this requires including additional internal variables in the PBE. The ODEs for $p_{b}$ and $m_{v}$ are initialized using an isothermal assumption and evolved at the quadrature nodes, keeping computation costs low. This strategy assumes their probability densities to be Dirac delta functions centered at the quadrature nodes for $R$ and $\dot{R}$. The ODEs for $p_{b}$ and $m_{v}$ follow Ando et al. [47] as

### v4

@@ 3.2.2 Approximate Riemann solver
The Riemann problem at each finite volume interface is solved using either the Harten–Lax–van Leer (HLL) or the Harten–Lax-van Leer contact (HLLC) approximate Riemann solver [Toro1997]. The HLL approximate Riemann solver admits three constant states separated by two discontinuities with different wave speeds. The wave speeds are estimated using the left and right state variables $\bm{q}^{\mathrm{(L)}}_{i+1/2,j,k}$ and $\bm{q}^{\mathrm{(R)}}_{i+1/2,j,k}$. The HLL flux satisfies the Rankine–Hugoniot conditions that ensure the conservation of mass, momentum, and energy across the discontinuity. A shortcoming of the HLL approximate Riemann solver is that it does not account for the contact discontinuity in the region between the two wave speeds (the so-called “star” region). As a result, the HLL approximate Riemann solver has more numerical dissipation than the HLLC approximate Riemann solver, which accounts for this third contact discontinuity. Continuity of normal velocity and pressure is imposed across the contact discontinuity, which enables computations of the star states and the resulting HLLC flux. Once computed, the HLL and HLLC fluxes are used to calculate the right-hand side of 10. Identical calculations are performed in the $x$- and $z$-directions.
@@ 3.3 Boundary conditions
Extrapolation and Characteristic [thompson1990] boundary conditions are implemented using buffer regions outside the domain to support the stencils of the finite volume reconstructions and provide information outside of the domain for characteristic boundary conditions. Subsonic and supersonic free-stream and wall boundary conditions are implemented to handle various simulation configurations. Simple boundary conditions remain a mainstay of MFC, which were established in bryngelson21. These boundary conditions include symmetry, slip and no-slip walls, prescription of inlet conditions along partial boundary regions, and first-order accurate extrapolation for Neumann-like boundary conditions.
@@ 3.4 Time integration
The conservative variables are advanced in time using high-order total variation diminishing (TVD) and strong stability preserving (SSP) explicit Runge–Kutta time steppers [Gottlieb1998]. First- and second-order accurate time-steppers Runge–Kutta are also available.
3.5 Convergence
The convergence of the existing numerical methods with the five- and six-equation models is verified using a one-dimensional two-component advection problem with periodic boundary conditions. The computational domain spans the interval $[0,1]$ with a uniform pressure and velocity equal to one. The volume fraction and density of each component are defined as $\alpha_{i}=\alpha_{i}\rho_{i}=0.5+\left(-1\right)^{i}\sin\left(2\pi x\right)$ and sum to one everywhere in the domain.
Figure 1: Convergence results for a 1D 2-component advection problem with the (a) HLL and (b) HLLC approximate Riemann solvers.
Figure 1 shows the RMS error after five periods of advection for $1^{st}$, $3^{rd}$, and $5^{th}$ order accurate reconstructions using the HLL and HLLC approximate Riemann solvers. Convergence results can be reproduced using the automated scripts provided in the examples/1D_convergence example case.
@@ 4 Updated capabilities
This section is organized by the role each additional physical, numerical, or performance capability plays in the overarching goal of providing an end-to-end state-of-the-art simulation platform for a broad range of flow problems. First, we broaden physical coverage from MFC 3.0 (see bryngelson21) to capture the couplings that govern realistic flows (immersed boundaries for complex geometry; N-fluid phase change; Euler–Euler and Euler–Lagrange bubble models; fluid–structure interaction; chemically reacting mixtures; two-material surface tension; MHD/RMHD). We equip the solver with the numerical machinery required to preserve accuracy and robustness across regimes (general characteristic boundary conditions, WENO-Z/TENO reconstructions, Strang splitting for stiff sub-grid dynamics, and low-Mach treatments). Lastly, we enhance the software’s performance and reproducibility, including portable GPU/APU offload, metaprogramming and code generation, continuous integration with over 500 regression tests, and extreme-scale I/O. So, these methods run consistently from a single accelerator to tens of thousands of devices.
@@ 4.1.1 Immersed boundary method
Fluid-solid interactions with rigid bodies are simulated using the Ghost-Cell Immersed Boundary Method with an implementation similar to tseng2003ghost. This method imposes Neumann boundary conditions for the pressure, densities, and volume fractions, as well as a no-slip or slip boundary condition for the velocity on the immersed boundary.
In MFC 5.0, the solid geometries are parsed. Each cell is characterized as being either within a fluid or a solid region. For each immersed boundary, the level set field $\bm{\varphi}$ is calculated. This field stores the vector from each cell to the closest point on the immersed boundary. The ghost points, image points, and interpolation coefficients are calculated and used to enforce the appropriate boundary conditions. These are computed before simulation execution and follow the algorithm below.
Algorithm 1 Calculation of level sets.
for each immersed boundary $\Gamma_{i}$ do
  for each cell do
   if $\Gamma_{i}$ is defined as a simple geometry then $\triangleright$ i.e: Sphere, Cuboid, etc.
     Calculate the level set with an analytical function.
   else if $\Gamma_{i}$ is defined by an STL/OBJ file then
     Calculate the level set using the procedure described in section 4.1.2.
   end if
  end for
end for
The algorithm that applies the boundary conditions follows. MFC first parses the STL or OBJ geometry file (or level set). We determine whether each finite volume is in a solid, immersed region by examining its surface normals. We use algorithm 1 to compute the level set field for each immersed boundary. This field stores the vector from each finite volume center to the closest point on the immersed boundary. If each cell center in the solid regions is within three cells of a fluid cell, we characterize it as a ghost point. The fluid properties at the image points are interpolated from surrounding grid cells. We follow algorithm 2 to compute the ghost points, image points, and the interpolation coefficients. Finite volumes in an immersed solid region are excluded from interpolation. Following algorithm 3, we compute $q_{\mathrm{(ip)}}=q_{\mathrm{(gp)}}$ for each fluid property $q$ with a Neumann boundary condition and set the velocity based on its boundary conditions.
@@
Algorithm 2 Determining Ghost Points, Image Points, and Interpolation Coefficients
for each cell do
  if the cell is in a solid region and it is within three cells of the fluid region then
   Add cell to the list of ghost points
  end if
end for
for each ghost point do
  $\bm{x}_{\mathrm{(ip)}}=\bm{x}_{\mathrm{(gp)}}+2\bm{\varphi}(\bm{x}_{\mathrm{(gp)}})$ $\triangleright$ $\bm{x}_{\mathrm{(ip)}}$ is the image point position, $\bm{x}_{\mathrm{(gp)}}$ is the ghost point position
  Determine the four grid cells surrounding the image point
  for each surrounding grid cell $\bm{g}_{j}(\bm{x}_{\mathrm{(ip)}})$ do
   if $\bm{g}_{j}(\bm{x}_{\mathrm{(ip)}})$ is in the fluid region then
     $d=\lVert\bm{x}_{\mathrm{(ip)}}-\bm{g}_{j}(\bm{x}_{\mathrm{(ip)}})\rVert_{2}$ $\triangleright$ Distance from image point to grid cell
     $c_{j}(\bm{x}_{\mathrm{(ip)}})=1/d^{2}$
   else
     $c_{j}(\bm{x}_{\mathrm{(ip)}})=0$
   end if
  end for
  $c(\bm{x}_{\mathrm{(ip)}})=c(\bm{x}_{\mathrm{(ip)}})/\sum_{j}c_{j}(\bm{x}_{\mathrm{(ip)}})$ $\triangleright$ Normalize the interpolation coefficients
end for
Algorithm 3 Ghost point treatment at each timestep.
for each ghost point do
  for each fluid property $q$ do
   $q(\bm{x}_{\mathrm{(ip)}})=\sum q(\bm{g}_{i}(\bm{x}_{\mathrm{(ip)}}))\cdot c_{i}(\bm{x}_{\mathrm{(ip)}})$
   if $q$ is velocity then
     if velocity boundary condition is no-slip then
      $\bm{u}(\bm{x}_{\mathrm{(gp)}})=0$
     else if velocity boundary condition is slip then
      $\bm{u}(\bm{x}_{\mathrm{(gp)}})=\bm{u}(\bm{x}_{\mathrm{(ip)}})-(\hat{\bm{n}}\cdot\bm{u}(\bm{x}_{\mathrm{(ip)}}))\hat{\bm{n}}$
     end if
[…]
     $q(\bm{x}_{\mathrm{(gp)}})=q(\bm{x}_{\mathrm{(ip)}})$
   end if
  end for
end for
@@ 4.1.2 Treatment of complex geometries
We use a ray-tracing algorithm to translate complex geometries in ASCII and stereolithography (STL) formats onto computational grids compatible with the immersed boundary method. The ASCII STL format represents geometric models using a triangular mesh. Each model surface is decomposed into a set of non-overlapping triangular facets. Three vertices and a normal vector define each triangular facet. Unlike simple geometries such as cubes and spheres, most complex geometric models do not have analytical solutions for their boundary normal vectors. Consequently, the level sets of complex geometries mentioned in section 4.1.1 also lack analytical solutions. We need to approximate the geometry’s boundary and boundary normal vectors using the STL vertices located on the boundary. We use an edge-manifoldness algorithm based on weiler1986topological to group the boundary STL vertices that are used to determine the level sets and image points. Coarse STL files are interpolated during ray tracing to achieve results consistent with those obtained from higher-resolution STL files. The implementation is validated for a Mach 2 helium flow over a sphere with initial flow density $\rho_{0}$ and $\mbox{\text{Re}}=6.5\times 10^{5}$ in fig. 2.
@@ 4.1.3 Phase change: First order transition, liquid–vapor
The phase change model for $N$ fluids is integrated into MFC implemented by introducing disequilibrium terms for pressure ($p$), temperature ($T$), and chemical potential ($g$) into $\bm{s}(\bm{q})$ in 9. This inclusion increases the system stiffness, making a fractional-step method essential [Strang1968, Zein2010, saurel2009]. In the case of infinitely fast relaxation, the governing PDEs can be replaced with nonlinear, lower-dimensional, algebraic relations. For example, for the $pT$-relaxation, we have [flaatten2011solutions]
@@
with $\alpha_{m}\rho_{m}=\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2}$, $e_{*,i}$ a reference parameter added to the stiffened gas EOS (see section 2.2) internal energy equation, and $c_{v,i}$ and $c_{p,i}$ the specific heat coefficient at constant volume and pressure, respectively, for species $i$. Subscripts $(\,\cdot\,)_{1,2}$ indicate the reacting liquid and vapor. The remaining subscripts are the inert phases. $A$, $B$, $C$, and $D$ are fluid-dependent parameters [saurel2008], reproduced here for clarity:
$A=\frac{c_{p,1}-c_{p,2}+e^{\prime}_{*,2}-e^{\prime}_{*,1}}{c_{p,2}-c_{v,2}},\quad B=\frac{e_{*,1}-e_{*,2}}{c_{p,2}-c_{v,2}},\quad C=\frac{c_{p,2}-c_{p,1}}{c_{p,2}-c_{v,2}},\quad D=\frac{c_{p,1}-c_{v,1}}{c_{p,2}-c_{v,2}},$
in which $e^{\prime}_{*,i}$ is a reference entropy parameter of the stiffened gas EOS for species $i$. These strategies are compatible with five- and six-equation models and are solved via Newton’s method. Once the solution is obtained, the solver advances to the next time step.
[…]
$U_{F}=\frac{(\rho u)_{a}-(\rho u)_{b}}{\rho_{a}-\rho_{b}}.$
Figure 3: Comparison between the numerical and experimental values for the evaporation front velocity using the 5- and 6-equation models for the chocked flow series of n-dodecane tests [simoesmoreira1999evaporation].
Figure 3 compares the numerical (5- and 6-equation models) and experimental results. We observe that both numerical results match well, with some minor discrepancies noted. These are suspected to be due to the sensitivity in the data obtained to calculate $U_{F}$, as the precise locations of “a” and “b” are unclear. The numerical results quantitatively match the experimental data, with closer matching at larger $T$.
@@ 4.1.4 Euler–Euler sub-grid bubble dynamics
We use Euler–Euler sub-grid models for bubble dynamics, including the method of classes and the method of moments. The method of classes is based on the ensemble phase-averaged equations [zhang1994], which can be used to represent dispersions of radially oscillating bubbles as well as other particles [bryngelson19_IJMF, bryngelson23, Bryngelson2020, bryngelson19_whales]. Note that the method of classes approach was included in MFC 3.0 [bryngelson21], though a statistical representation enabled via the moment method was not included, which we focus on here.
The void fraction of the dispersed phase $\alpha$ is assumed to be negligible compared to the liquid phase ($\alpha_{l}$) under dilute assumptions. The bubbles are represented statistically via random variables $R$, $\dot{R}$, and $R_{o}$, corresponding to the instantaneous bubble radius, time derivative, and equilibrium bubble radius. The mixture-averaged pressure field is
@@
where $p_{bw}$ is the associated bubble wall pressure and $p_{\ell}(\bm{x},t)$ is the liquid pressure according to the modified stiffened-gas EOS [menikoff89]:
@@
The right-hand side represents the change in void fraction associated with bubble growth and collapse. The over-barred terms appearing in the above equations denote average quantities of the bubble dispersion. These are averaged over $N_{\mathrm{bin}}$ bins of $R_{o}$ using a log-normal probability distribution function (pdf). Hybrid schemes and fast integration techniques have been developed for polydisperse cases that otherwise require large bin counts [charalampopoulos21, bryngelson_levin26]. We implement the Rayleigh–Plesset and Keller–Miksis equations [plesset77] for the radial bubble dynamics. The Keller–Miksis dynamics are
@@
The method of classes incorporates stochasticity in $R_{o}$ using a log-normal probability density function (pdf). However, modeling complex bubbly flow phenomena requires the expansion of the set of stochastic variables, including instantaneous bubble variables. The inaccuracy of the polytropic assumption can also limit its scope of application. The method of moments is employed for this purpose, utilizing a population balance equation (PBE) to govern the probability density function (pdf) of the bubbles. The pdf $f(R,\dot{R},R_{o})$ governs the moments $\mu_{lmn}$ as
@@
The moments are evaluated using a conditional inversion procedure that follows Fox2018. The conditional probability density function $f(R,\dot{R}|R_{o})$ without coalescence or breakup effects is governed by the PBE
@@
and $\Omega=\Omega_{R}\times\Omega_{\dot{R}}\times\Omega_{R_{o}}=(0,\infty)\times(-\infty,\infty)\times(0,\infty)$ [Bryngelson2020]. The integral is computed via quadrature nodes obtained by the inversion procedure.
@@
where $N_{R_{o}}$ represents the polydispersity of the bubble population. The conditional moments $\mu_{lm}$ are evaluated using a conditional hyperbolic inversion procedure (CHyQMOM). The CHyQMOM algorithm is described in Fox2018, and a detailed discussion of its application to bubble dynamics can be found in bryngelson23. The total moments follow as
@@
These moments evaluate the right-hand side of the moment evolution equation and the ensemble-phase averaged terms. The evaluation of $\overline{R^{3}p_{bw}}$ requires the accounting of statistics of $p_{b}$. Without polytropic assumptions, this requires including additional internal variables in the PBE. The ODEs for $p_{b}$ and $m_{v}$ are initialized using an isothermal assumption and evolved at the quadrature nodes, keeping computation costs low. This strategy assumes their probability densities to be Dirac delta functions centered at the quadrature nodes for $R$ and $\dot{R}$. The ODEs for $p_{b}$ and $m_{v}$ follow ando09 as